# Constrained Optimization From a Control Perspective via Feedback Linearization

### v2

@@ Abstract
Tools from control and dynamical systems have proven valuable for analyzing and developing optimization methods. In this paper, we establish rigorous theoretical foundations for using feedback linearization—a well-established nonlinear control technique—to solve constrained optimization problems. For equality-constrained optimization, we establish global convergence rates to first-order Karush-Kuhn-Tucker (KKT) points and uncover the close connection between the FL method and the Sequential Quadratic Programming (SQP) algorithm. Building on this relationship, we extend the FL approach to handle inequality-constrained problems. Furthermore, we introduce a momentum-accelerated feedback linearization algorithm that achieves faster convergence, and provides a rigorous convergence guarantee.
@@ 1 Introduction
Recent studies Cerone et al. (2024); Gunjal et al. (2024) have explored the dynamical properties of continuous time constrained optimization algorithms. These works leverage a feedback control perspective to design and analyze the performance of optimization methods. Specifically, they propose frameworks that model constrained optimization problems as control problems, where the iterations of the optimization algorithm are represented by a dynamical system, and the Lagrange multipliers act as control inputs. The objective in this framework is to drive the system to a feasible steady state that satisfies the constraints. Within this framework, various control strategies can be employed to design the update of Lagrange multipliers, resulting in different control-based first-order methods.
In this work, we adopt the same control perspective as above, and specifically focus on using Feedback Linearization (FL) approach, a standard approach in nonlinear control (cf. Isidori (1985); Henson and Seborg (1997)), to design the Lagrange multiplier. One key advantage of this method is its natural suitability for handling nonconvex constrained optimization problems. Although this approach has been explored in the literature Cerone et al. (2024); Schropp and Singer (2000), its theoretical properties are not yet fully understood. Several important questions remain open.
The first question concerns global convergence and convergence rates. While existing works established local stability Cerone et al. (2024), global convergence and convergence rate have not been established. The second question concerns the relationship between the feedback linearization approach and existing optimization algorithms, specifically whether the discretization of the optimization dynamics derived from feedback linearization aligns with any known optimization method. Additionally, since most existing studies Cerone et al. (2024); Schropp and Singer (2000) focus exclusively on equality constraints, this raises the third question: how can the feedback linearization approach be extended to effectively handle inequality constraints? Lastly, it remains an open question whether insights from acceleration techniques in optimization (e.g., momentum acceleration) can be leveraged to develop faster feedback linearization algorithms, leading to faster algorithms for constrained optimization.
@@
    Finally, leveraging these findings, we propose a momentum-accelerated FL algorithm for constrained optimization, which empirically achieves accelerated convergence. Additionally, we provide a rigorous convergence guarantee for the continuous-time momentum-accelerated FL method (Section 5). To the best of our knowledge, both the proposed algorithm and its analysis are novel contributions to constrained optimization.
@@
We use the notation $n],n\in\mathbb{N}$ to denote the set $\{1,2,3,\dots,n\}$. We use $\nabla f(x)$ to denote the gradient of a scalar function $f:\mathbb{R}^{n}\to\mathbb{R}$ evaluated at the point $x\in\mathbb{R}^{n}$ and use $\nabla^{2}f(x)$ to denote its corresponding Hessian matrix. We use $J_{h}(x)$ to denote the Jacobian matrix of a function $h:\mathbb{R}^{n}\to\mathbb{R}^{m}$ evaluated at $x\in\mathbb{R}^{n}$, i.e. $J_{h}(x)]_{i,j}=\frac{\partial h_{i}(x)}{\partial x_{j}}$, $i\in[m],j\in[n]$. Unless specified otherwise, we use $\|\cdot\|$ to denote the $L_{2}$ norm of matrices and vectors and use $\|\cdot\|_{\infty}$ to denote the $L_{\infty}$ norm. For a positive definite matrix $A$, we use $\|X\|_{A}:=\|A^{-\frac{1}{2}}X\|$ to denote the $A$-norm of $X$. For a set $\mathcal{A}$, we use $\mathcal{A}^{c}$ to denote its complement.
@@ 2 Feedback Linearization (FL) for solving equality constrained optimization
where $x\in\mathbb{R}^{n}$, $f:\mathbb{R}^{n}\to\mathbb{R},h:\mathbb{R}^{n}\to\mathbb{R}^{m}$. The first-order KKT conditions are given by
$\begin{split}-\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)=0\end{split}$ (2)
@@
$\begin{split}&\frac{dx}{dt}=-T(x(t))\left(\nabla f(x(t))+J_{h}(x(t))^{\top}\lambda(t)\right),\\ &y(t)=h(x(t)),\end{split}$ (3)
@@
${\lambda_{\min}}I\preceq T(x)\preceq\lambda_{\max}I$
For notational simplicity, in the following sections of the paper, we denote $\frac{dx}{dt}$ as $\dot{x}$ and abbreviate $x(t),\lambda(t),y(t)$ as $x,\lambda,y$.
Note that at an equilibrium point $x^{\star}$ of the system in Fig. 1 it must satisfy:
$\textstyle\dot{x}=0~\Longrightarrow~\nabla f(x^{\star})+J_{h}^{\top}(x^{\star})=0.$
Further, if $x^{\star}$ is feasible, i.e. $h(x^{\star})=0$, then we get that $x^{\star}$ satisfies the first order KKT conditions (2). Thus, the key idea is to manipulate the evolution of $x$ so that we stabilize the system to equilibrium and feasibility. (For a more detailed overview about optimization from a control perspective, see Appendix A.)
@@
$\dot{y}=J_{h}(x)\dot{x}=\underbrace{-J_{h}(x)T(x)}_{F(x)}\nabla f(x)\underbrace{-{J_{h}(x)T(x)J_{h}(x)^{\top}}}_{G(x)}\lambda$
@@
$\lambda=-\!\left(J_{h}(x)T(x)J_{h}(x)^{\!\top}\!\right)^{\!-\!1}\!\!(u+J_{h}(x)T(x)\nabla f(x))$
we have that $\dot{y}=u$, then we can simply set u=Kyu=-Ky where $K$ is a Hurwitz matrix to guarantee that $y$ asymptotically converge to zero. Thus the feedback linearization (FL) dynamics is given as follows:
FL for Equality-Constrained Optimization Cerone et al. (2024) $\vskip-20.0pt\dot{x}$ $=-T(x)\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)$ (5) $\lambda$ $\!=\!-\!\left(J_{h}(x)T(x)J_{h}(x)^{\!\top}\!\right)^{\!-\!1}\!\!(J_{h}(x)T(x)\nabla f(x)\!-\!Kh(x))$
The feedback linearization (FL) approach is particularly effective for handling nonlinear dynamics, making it well-suited for nonconvex constrained optimization. Numerical results in Cerone et al. (2024); Schropp and Singer (2000) highlight its strong performance in such settings. However, its theoretical properties remain less well understood. Existing analyses primarily focus on local stability Cerone et al. (2024), while global convergence and convergence rates are largely unexplored. Additionally, the connection between the FL algorithm and existing optimization methods is not well established. It also remains unclear how to leverage the FL approach to develop novel techniques for inequality-constrained optimization and faster constrained optimization. In the following sections, we will systematically address these open problems.
@@ 3.1 Convergence Results
The result relies on the following assumption:
@@
Thefunctions $f$ and $h$ are assumed to satisfy the following conditions:
  1. 1.1
    There exists a constant D such that $(J_{h}(x)J_{h}(x)^{\top})^{-1}\prec D^{2}I$ for all $x$;
  2. 1.2
@@
Note that Assumption 1.1 is similar to the assumption made in Cerone et al. (2024) which assumes that $\textup{rank}(J_{h}(x))=m$ for all $x$, which is equivalent to $J_{h}(x)J_{h}(x)^{\top}$ is invertible. This assumption is also known as the linear independence constraint qualification (LICQ, cf. Peterson (1973); Nocedal and Wright (2006), see more discussion in Appendix A) in optimization literature. Assumption 1.2 implies that the functions $f$ and $g$ are Lipschitz. We would like to acknowledge that this Assumption is quite restrictive and is solely for analysis purpose. In our numerical simulations we found that the algorithm is suitable for non-uniformly-Lipschitz functions. 111 We note that if $x^{\star}\in D$ is known a priori for a compact domain $D$, a potential approach for handling non-uniformly Lipschitz functions $f,g$ is to construct smooth and Lipschitz extensions $f^{\prime},g^{\prime}$ such that their gradients and Jacobians match those of $f,g$ within $D$ while remaining uniformly Lipschitz outside $D$ (cf. Stein (1970)).
We define the KKT-gap of $(x,\lambda)$ as follows:
$\textup{{KKT-gap}}(x,\lambda)\!:=\!\max\{\|\nabla f(x)\!+\!J_{h}(x)^{\top}\lambda\|,\|h(x)\|_{\infty}\}$ (6)
We now state our result in terms of the convergence rate
@@
Under Assumption 1, for control gain $K$ that is a diagonal positive definite matrix, i.e., $K=\textup{diag}\{k_{i}\}_{i=1}^{m}$, where $k_{i}>0$, we have that the dynamic of feedback linearization method (5) satisfies:
@@
    For the set $\mathcal{E}_{i}:=\{x:h_{i}(x)\geq 0\}$, if $x(0)\in\mathcal{E}_{i}$, then $x(t)\in\mathcal{E}_{i}$ for all $t\geq 0$. Similarly, if $x(0)\in\mathcal{E}_{i}^{c}$, then $x(t)\in\mathcal{E}_{i}^{c}$ for all $t\geq 0$, further
    $h_{i}(x(t))=e^{-k_{i}t}h_{i}(x(0)),$
    i.e., $h(x(t))\to 0$ with an exponential rate as $t\to+\infty$.
@@
    Define $\ell(x)\!:=\!f(x)\!+\!\frac{\lambda_{\max}}{{\lambda_{\min}}}(MD)^{2}\sum_{i=1}^{m}|h_{i}(x)|$, then $\ell(x(t))$ is non-increasing with respect to $t$.
@@
    $\!\!\!\!\!\!\!\int_{t\!=\!0}^{T}\!\!\!\!\!\|\nabla\!f(\!x(t))\!+\!J_{h}(\!x(t))^{\!\top}\overline{\lambda}(t)\|^{2}dt\!\leq\!\!\frac{1}{{\lambda_{\min}}}\!\left(\ell(x(0))\!-\!\ell(x(T))\right),$
@@
    $\quad\inf_{0\leq t\leq T}\textup{{KKT-gap}}(x(t),\overline{\lambda}(t))$
    $\leq\max\!\left\{\!\!\sqrt{\frac{2}{T}\!\left(\frac{f(x(0))\!-\!f_{\min}}{{\lambda_{\min}}}\!+\!\frac{\lambda_{\max}M^{2}D^{2}}{{\lambda_{\min}}^{2}}\sum_{i}|h_{i}(x(0))|\right)},\max_{1\leq i\leq m}\left\{h_{i}(x(0))e^{-\frac{k_{i}T}{2}}\right\}\right\}\sim O\left(\frac{1}{\sqrt{T}}\right)$
    further, we have that
Statement 4 in Theorem 1 implies that the algorithm can find an $\epsilon$-first-order-KKT-point within time $\frac{1}{\epsilon^{2}}$. We note that ensuring last-iterate convergence in nonconvex optimization is generally challenging. Hence, our analysis focuses on the best iterate, a widely adopted criterion in nonconvex optimization. Due to space limitations, we defer the detailed proof in Appendix B. The key step of the proof involves constructing the Lyapunov function $\ell(x)$ in Statement 2. We would also like to note that $\ell(x)$ also serves as the exact penalty function in constrained optimization literature (cf. Eremin (1967); Zangwill (1967)).

### v3

@@ Abstract
Tools from control and dynamical systems have proven valuable for analyzing and developing optimization methods. In this paper, we establish rigorous theoretical foundations for using feedback linearization (FL)—a well-established nonlinear control technique—to solve constrained optimization problems. For equality-constrained optimization, we establish global convergence rates to first-order Karush-Kuhn-Tucker (KKT) points and uncover the close connection between the FL method and the Sequential Quadratic Programming (SQP) algorithm. Building on this relationship, we extend the FL approach to handle inequality-constrained problems. Furthermore, we introduce a momentum-accelerated feedback linearization algorithm and provide a rigorous convergence guarantee.
@@ 1 Introduction
Recent studies (c.f. Cerone et al. (2025); Gunjal et al. (2024); Allibhoy and Cortés (2024); Muehlebach and Jordan (2022a); Chen et al. (2023, 2022); França et al. (2018)) have explored the dynamical properties of continuous time constrained optimization algorithms. These works leverage a feedback control perspective to design and analyze the performance of optimization methods. Specifically, they propose frameworks that model constrained optimization problems as control problems, where the iterations of the optimization algorithm are represented by a dynamical system, and the Lagrange multipliers act as control inputs. The objective in this framework is to drive the system to a feasible steady state that satisfies the constraints. Within this framework, various control strategies can be employed to design the update of Lagrange multipliers, resulting in different control-based first-order methods. For example, it can be shown that Proportional-Integral (PI) control leads to Primal-Dual Gradient Dynamics (PDGD), whose properties are well-studied Kose (1956); Qu and Li (2018); Ding and Jovanović (2019). However, most of the works focuses on convergence for convex constrained problems.
In this work, we adopt the same control perspective as above, and specifically focus on using another approach, namely Feedback Linearization (FL) a standard approach in nonlinear control (cf. Isidori (1985); Henson and Seborg (1997)), to design the Lagrange multiplier. One key advantage of this method is its natural suitability for handling nonconvex constrained optimization problems. Although this approach Cerone et al. (2025), along with similar dynamical system perspectives Schropp and Singer (2000); Allibhoy and Cortés (2024); Muehlebach and Jordan (2022b, 2023), has been explored in the literature, its theoretical properties are not yet fully understood. Several important questions remain open.
The first question concerns global convergence and convergence rates. While existing works established local stability Cerone et al. (2025), global convergence and convergence rate have not been established. The second question concerns the relationship between the feedback linearization approach and existing optimization algorithms, specifically whether the discretization of the optimization dynamics derived from feedback linearization aligns with any known optimization method. Additionally, since most existing studies Cerone et al. (2025); Schropp and Singer (2000) focus exclusively on equality constraints, this raises the third question: how can the feedback linearization approach be extended effectively for inequality constraints? Lastly, it remains unclear whether ideas from acceleration in optimization–such as momentum-based techniques–can be incorporated to speed up feedback lienarization methods for constrained optimization.
@@
    Finally, we propose a momentum-accelerated FL algorithm for constrained optimization, which empirically exhibits accelerated convergence in both equality constrained and inequality constrained settings. Furthermore, we establish $O(\frac{1}{\sqrt{T}})$ convergence in the nonconvex equality constrained setting (Section 5).
@@
We use the notation $n],n\in\mathbb{N}$ to denote the set $\{1,2,3,\dots,n\}$. We use $\nabla f(x)$ to denote the gradient of a scalar function $f:\mathbb{R}^{n}\to\mathbb{R}$ evaluated at the point $x\in\mathbb{R}^{n}$ and use $\nabla^{2}f(x)$ to denote its corresponding Hessian matrix. We use $J_{h}(x)$ to denote the Jacobian matrix of a function $h:\mathbb{R}^{n}\to\mathbb{R}^{m}$ evaluated at $x\in\mathbb{R}^{n}$, i.e. $J_{h}(x)]_{i,j}=\frac{\partial h_{i}(x)}{\partial x_{j}}$, $i\in[m],j\in[n]$. Unless specified otherwise, we use $\|\cdot\|$ to denote the $L_{2}$ norm of matrices and vectors and use $\|\cdot\|_{\infty}$ to denote the $L_{\infty}$ norm. For a positive definite matrix $A$, we use $\|X\|_{A}:=\|A^{-\frac{1}{2}}X\|$ to denote the $A$-norm of $X$. For a set $\mathcal{A}$, we use $\mathcal{A}^{c}$ to denote its complement. When no ambiguity arises, we denote $\frac{dx}{dt}$ by $\dot{x}$, and abbreviate time-dependent variables such as $x(t)$, $\lambda(t)$, and $y(t)$ as $x$, $\lambda$, and $y$.
@@ 2 Feedback Linearization (FL) for solving equality constrained optimization
where $x\in\mathbb{R}^{n}$, $f:\mathbb{R}^{n}\to\mathbb{R},h:\mathbb{R}^{n}\to\mathbb{R}^{m}$. Here we assume that $f,h$ are differentiable, and additional assumptions will be introduced where needed to support the analysis. The first-order KKT conditions are given by
$\begin{split}\textstyle-\nabla f(x)-J_{h}(x)^{\top}\lambda=0,\quad h(x)=0\end{split}$ (2)
@@
$\begin{split}&\textstyle\frac{dx}{dt}=-T(x(t))\left(\nabla f(x(t))+J_{h}(x(t))^{\top}\lambda(t)\right),\\ &\textstyle y(t)=h(x(t)),\end{split}$ (3)
$\textstyle{\lambda_{\min}}I\preceq T(x)\preceq\lambda_{\max}I$
@@
Note that at an equilibrium point $x^{\star}$ of the system in Fig. 1 must satisfy: $\dot{x}=0~\Longrightarrow~\nabla f(x^{\star})+J_{h}^{\top}(x^{\star})\lambda=0.$
Further, if $x^{\star}$ is feasible, i.e. $h(x^{\star})=0$, then we get that $x^{\star}$ satisfies the first order KKT conditions (2). Thus, the key idea is to manipulate the evolution of $x$ so that we stabilize the system to equilibrium and feasibility.
To measure convergence, We define the KKT-gap of $(x,\lambda)$ as follows, such that converging to a KKT point is equivalent to KKT-gap is zero  [^1] In this paper, we use the term ‘convergence’ to refer to the decay of the KKT gap to zero. While this is weaker than convergence to a local or global optimum, it remains a meaningful guarantee in nonconvex constrained optimization. Extending the framework to incorporate saddle-point escape techniques and ensure convergence to local optima is an important direction for future work.:
$\textstyle\textup{{KKT-gap}}(x,\lambda):=\max\{\|\nabla f(x)+J_{h}(x)^{\top}\lambda\|,\|h(x)\|_{\infty}\}.$ (4)
@@
$\textstyle\dot{y}=J_{h}(x)\dot{x}=\underbrace{-J_{h}(x)T(x)}_{F(x)}\nabla f(x)\underbrace{-{J_{h}(x)T(x)J_{h}(x)^{\top}}}_{G(x)}\lambda$
@@
$\textstyle\lambda=-\left(J_{h}(x)T(x)J_{h}(x)^{\top}\right)^{-1}(u+J_{h}(x)T(x)\nabla f(x))$
we have that $\dot{y}=u$, then we can simply set u=Kyu=-Ky where $K$ is a Hurwitz matrix to guarantee that $y$ asymptotically converge to zero. Thus the feedback linearization (FL) dynamics is given as:
FL for Equality-Constrained Optimization Cerone et al. (2025) $\begin{split}\dot{x}&=-T(x)\left(\nabla f(x)+J_{h}(x)^{\top}\lambda\right)\\ \lambda&=-\left(J_{h}(x)T(x)J_{h}(x)^{\top}\right)^{-1}(J_{h}(x)T(x)\nabla f(x)-Kh(x))\end{split}$ (6)
The FL approach is particularly effective for handling nonlinear dynamics, making it well-suited for nonconvex constrained optimization. Numerical results in Cerone et al. (2025); Schropp and Singer (2000) highlight its strong performance in such settings. However, its theoretical properties remain less well understood. Existing analyses primarily focus on local stability Cerone et al. (2025), while global convergence and convergence rates are largely unexplored. Additionally, the connection between the FL algorithm and existing optimization methods is not well established. It also remains unclear how to leverage the FL approach to develop novel techniques for inequality-constrained optimization and faster constrained optimization. In the following sections, we will systematically address these open problems.
Remark 1 (Scalability and Computational Complexity).
Note that although (6) requires calculating the matrix inversion of ${J_{h}(x)T(x)J_{h}(x)^{\top}}\in\mathbb{R}^{m\times m}$, the dimension scales with the number of constraints $m$ instead of the dimension of the optimization variable $x\in\mathbb{R}^{n}$. In many practical settings, e.g. safe RL, the number of constraints is significantly smaller than the dimension of $x$. In such cases, the inversion is computationally inexpensive and can be performed efficiently. Moreover, even in cases where the number of constraints is large and the matrix inversion becomes computationally burdensome, we show that it is possible to approximate the inverse efficiently while still maintaining convergence guarantees. Specifically, in Appendix B.2, we present a modified FL algorithm that incorporates a Proportional-Integral (PI) controller to tolerate approximation error and still ensure convergence to an optimal solution.
Remark 2 (Extensions of the FL approach).
Although similar dynamics to (6) has also been proposed from other perspectives such as control barrier functions Allibhoy and Cortés (2024), nonsmooth dynamics design Muehlebach and Jordan (2022a, 2024), we believe that the feedback linearization perspective provides a more general framework. In principle, beyond specifying a linear target of the form $\dot{y}=-Ky$, this approach allows us to leverage arbitrary stable controllers, such as PI controllers, e.g.
$\textstyle\dot{y}=-K_{p}y-K_{i}\int_{0}^{t}y(s)ds$ (7)
for constraint enforcement, potentially offering new algorithmic behaviors or robustness benefits. In Appendix B.2, we present an example where the algorithm in (7) succeeds in converging to the optimal solution, even when only an inaccurate approximation of $\left({J_{h}(x)T(x)J_{h}(x)^{\top}}\right)^{-1}$ is available—while (6) fails to converge to an optimal solution. This highlights the robustness and broader applicability of the feedback linearization framework.
@@ 3.1 Convergence Results
The result relies on the following assumptions:
@@
Assumption 3.
There exists a constant D such that $(J_{h}(x)J_{h}(x)^{\top})^{-1}\prec D^{2}I$ for all $x$.
Note that Assumption 3 is similar to the assumption made in Cerone et al. (2025) that assumes that $\textup{rank}(J_{h}(x))=m$ for all $x$, which is equivalent to $J_{h}(x)J_{h}(x)^{\top}$ being invertible, thereby ensuring the regularity of the transformation in (6) and the existence of a well-defined feedback linearization. This assumption is also known as the linear independence constraint qualification (LICQ, cf. Peterson (1973); Nocedal and Wright (2006), see more discussion in Appendix A) in optimization literature. Assumption 1 implies that the functions $f$ and $g$ are Lipschitz. We would like to acknowledge that this assumption is relatively restrictive and is solely for analysis purpose  [^2] We note that if $x^{\star}\in D$ is known a priori for a compact domain $D$, a potential approach for handling non-uniformly Lipschitz functions $f,g$ is to construct Lipschitz extensions $f^{\prime},g^{\prime}$ such that their gradients and Jacobians match those of $f,g$ within $D$ while remaining uniformly Lipschitz outside $D$ (cf. Stein (1970)).. In our numerical simulations we found that the algorithm is suitable for non-uniformly-Lipschitz functions. We now state our result in terms of the convergence rate:
@@
Let Assumption 1, 2 and 3 hold and let the control gain $K$ be a diagonal positive definite matrix, i.e., $K=\textup{diag}\{k_{i}\}_{i=1}^{m}$, where $k_{i}>0$. Then we have that the dynamic of the feedback linearization method (6) satisfies:
@@
    For the set $\mathcal{E}_{i}:=\{x:h_{i}(x)\geq 0\}$, if $x(0)\in\mathcal{E}_{i}$, then $x(t)\in\mathcal{E}_{i}$ for all $t\geq 0$. Similarly, if $x(0)\in\mathcal{E}_{i}^{c}$, then $x(t)\in\mathcal{E}_{i}^{c}$ for all $t\geq 0$, further $h_{i}(x(t))=e^{-k_{i}t}h_{i}(x(0))$, i.e., $h(x(t))\to 0$ with an exponential rate as $t\to+\infty$.
@@
    Define $\ell(x):=f(x)+\frac{\lambda_{\max}}{{\lambda_{\min}}}(MD)^{2}\sum_{i=1}^{m}|h_{i}(x)|$, then $\ell(x(t))$ is non-increasing w.r.t. $t$.
@@
    $\textstyle\int_{t=0}^{T}\|\nabla f(x(t))+J_{h}(x(t))^{\top}\overline{\lambda}(t)\|^{2}dt\leq\frac{1}{{\lambda_{\min}}}\left(\ell(x(0))-\ell(x(T))\right),$
@@
    $\textstyle\inf_{0\leq t\leq T}\textup{{KKT-gap}}(x(t),\overline{\lambda}(t))\leq$ $\textstyle\max\left\{\sqrt{\frac{2}{T}\left(\frac{f(x(0))-f_{\min}}{{\lambda_{\min}}}+\frac{\lambda_{\max}M^{2}D^{2}}{{\lambda_{\min}}^{2}}\sum_{i}|h_{i}(x(0))|\right)},\right$.
    $\textstyle\hskip 18.49988pt\hskip 18.49988pt\left.\max_{1\leq i\leq m}\left\{h_{i}(x(0))e^{-\frac{k_{i}T}{2}}\right\}\right\}\sim O\left(\frac{1}{\sqrt{T}}\right)$
    further, we have that $\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\overline{\lambda}(t))=0,~\lim_{t\to+\infty}\textup{{KKT-gap}}(x(t),\lambda(t))=0.$
Statement 4 in Theorem 1 implies that the algorithm can find an $\epsilon$-first-order-KKT-point within time $\frac{1}{\epsilon^{2}}$. We note that ensuring last-iterate convergence in nonconvex optimization is generally challenging. Hence, our analysis focuses on the best iterate, a widely adopted criterion in nonconvex optimization. However, in the setting where the optimization problem (1) is strongly convex, we are able to strengthen our convergence result to the last iterate convergence to the global optimal solution. Due to space limitations, we defer the detailed proof of Theorem 1 as well as the result for the strongly convex setting to Appendix C. The key step of the proof involves constructing the merit function $\ell(x)$ in Statement 2. We also note that $\ell(x)$ also serves as the exact penalty function in constrained optimization literature (cf. Eremin (1967); Zangwill (1967)).